OpenStudy (howard-wolowitz):

The first three books in the series have 117, 135, and 117 pages. The author finishes the series with a fourth book that has 151 pages. Which is not true? A. The range is smaller than the mean. B. The mean is smaller than the median. C. The mean is larger than the mode. D. The median is larger than the mode.

OpenStudy (solomonzelman):

The range is maximum-minimum.

OpenStudy (howard-wolowitz):

Yes! Were on the same page

OpenStudy (solomonzelman):

I am sure you can tell that max−min=151−117<100 And mean>100 since all values are greater than 100.

OpenStudy (howard-wolowitz):

ok so that means the mean is smaller then the rest

OpenStudy (howard-wolowitz):

so I think the answer is B

OpenStudy (howard-wolowitz):

the median of all the rest of the values, is ... the median

OpenStudy (solomonzelman):

Median = (117+135)/2 Mean = (117+135+117+151)/ =(117+135+117+135+positive)/4 =(117+135+117+135)/4 + positive/4 =(2•(117+135))/4 + positive =(117+135)/2 + positive

OpenStudy (solomonzelman):

I proved that mean is larger than the median

OpenStudy (solomonzelman):

So B is wrong... see how?

OpenStudy (howard-wolowitz):

yes but thats upsetting

OpenStudy (solomonzelman):

I gave you the hint

OpenStudy (solomonzelman):

max−min=151−117<100 right ?

OpenStudy (howard-wolowitz):

yes right I tihnk

OpenStudy (solomonzelman):

good

OpenStudy (solomonzelman):

If all values are greater than 100, then their mean is also greater than 100

OpenStudy (solomonzelman):

is that reasonable?

OpenStudy (solomonzelman):

Oh, the question is not true! I am sorry!! Then, B is right

OpenStudy (solomonzelman):

but, let's still work other options, if you don't mind.

OpenStudy (howard-wolowitz):

ok

OpenStudy (solomonzelman):

Ok, so Range = max - min = 151−117<100 Mean > 100 (since all values are greater than 100)

OpenStudy (solomonzelman):

Which option is true based on that?

OpenStudy (howard-wolowitz):

mode

OpenStudy (howard-wolowitz):

are you asking me whats bigger then the mean?

OpenStudy (solomonzelman):

from my post, Range<100, Mean>100. So, Range<Mean.

OpenStudy (solomonzelman):

So based on that statement A is true.

OpenStudy (solomonzelman):

117, 135, 117, 151

OpenStudy (solomonzelman):

Is mean greater or smaller than 117?

OpenStudy (howard-wolowitz):

you said mean is Mean>100

OpenStudy (solomonzelman):

The smallest value is 117, so how can the mean be smaller than 100?

OpenStudy (howard-wolowitz):

i didnt say it was

OpenStudy (solomonzelman):

well, regardless of who "they" is, let's ignore them.

OpenStudy (howard-wolowitz):

well its my fault i confused myself

OpenStudy (solomonzelman):

Ok

OpenStudy (solomonzelman):

So, we know that A is true, right? (i've shown that before)

OpenStudy (howard-wolowitz):

yes

OpenStudy (solomonzelman):

The mode is 117. (most frequent value) The smallest value is 117, and some values are greater than 117. Therefore the mean is greater than 117.

OpenStudy (solomonzelman):

So if, mode=117 mean>117 then which statement is (also) true?